# On Large-scale Evaluation of Embedding Models for Knowledge Graph Completion

### v1

@@ Abstract
Knowledge graph embedding (KGE) models are extensively studied for knowledge graph completion, yet their evaluation remains constrained by unrealistic benchmarks. Commonly used datasets are either faulty or too small to reflect real-world data. Few studies examine the role of mediator nodes, which are essential for modeling n-ary relationships, or investigate model performance variation across domains. Standard evaluation metrics rely on the closed-world assumption, which penalizes models for correctly predicting missing triples, contradicting the fundamental goals of link prediction. These metrics often compress accuracy assessment into a single value, obscuring models’ specific strengths and weaknesses. The prevailing evaluation protocol operates under the unrealistic assumption that an entity’s properties, for which values are to be predicted, are known in advance. While alternative protocols such as property prediction, entity-pair ranking and triple classification address some of these limitations, they remain underutilized. This paper conducts a comprehensive evaluation of four representative KGE models on large-scale datasets FB-CVT-REV and FB+CVT-REV. Our analysis reveals critical insights, including substantial performance variations between small and large datasets, both in relative rankings and absolute metrics, systematic overestimation of model capabilities when n-ary relations are binarized, and fundamental limitations in current evaluation protocols and metrics.
@@ 1 Introduction
In this paper, we present a comprehensive large-scale evaluation of KGE models for KG completion. Although these models have been extensively evaluated to assess their efficacy and applicability in real-world applications, we demonstrate fundamental limitations in current evaluation methodologies that result in inaccurate assessments of model performance. Our investigation focuses particularly on large-scale analysis, examining three interrelated issues: the influence of data representation choices on performance outcomes, inherent weaknesses in evaluation protocols, and shortcomings of standard rank-based evaluation metrics. Our findings reveal that these methodological limitations substantially affect how KGE models are understood and compared within the research community.
@@
Many existing evaluations of KGE models rely on datasets, such as FB15k-237 [6] and WN18RR [9], both of which exhibit several limitations. First, they fail to capture n-ary relationships among entities, which are modeled using mediator nodes (also called CVT nodes) in the Freebase data dump [25] and statement nodes in Wikidata.11todo: 1Our earlier… Our earlier findings [34] indicate that the conversion of n-ary relationships to binary ones, as implemented in FB15k-237, leads to substantial overestimation of KGE models’ performance. Second, these datasets are considerably smaller than real-world datasets. Experiment results from previous studies [34, 47] reveal significant differences in the performance of models between small-scale and large-scale datasets, with relative performance rankings among models also varying based on dataset size. In this paper, we perform our experiments on two large-scale datasets, FB-CVT-REV and FB+CVT-REV [34]. While FB+CVT-REV represents n-ary relationships using CVT nodes, FB-CVT-REV converts n-ary relationships into binary ones, similar to FB15k-237. This distinction allows us to assess how KGE model performance varies when data is modeled differently. These datasets, derived from Freebase, are large-scale, represent real-world scenarios, and do not include the data redundancy that exists in FB15k. A discussion of these two datasets and other benchmarks is provided in Section 2.
@@
KG completion methods aim to identify missing but correct triples in a KG. However, current evaluation metrics paradoxically penalize models for correctly predicting triples that are missing from the evaluation dataset. Standard ranking-based metrics such as Mean Rank (MR) and Mean Reciprocal Rank (MRR) [45] operate under the closed-world assumption, treating all predictions outside the KG as incorrect. This contradicts the open-world assumption, which acknowledges that absent facts are not necessarily false [28]. Furthermore, each of these metrics aggregates performance across all relations and triples into a single value, obscuring models’ strengths and weaknesses on different relations. In Section 3, we explore the shortcomings of existing evaluation metrics with supporting experimental evidence. We illustrate how these shortcomings impact models’ performance and propose refinements to improve evaluation reliability.
@@
Besides the limitations of commonly used benchmark datasets and evaluation metrics, the standard evaluation protocol for KGE models, link prediction [6], also has shortcomings. Link prediction evaluates models by predicting the missing h in triple (?, r, t) or missing t in (h, r, ?). It assesses whether a model prioritizes correct answers over incorrect ones. However, as Wang et al. [40] pointed out, this approach can be misleading because it does not verify whether correct triples are ranked better than all type of false and nonsensical ones. For example, for a test triple (James Ivory, director/film, A Room With A View), evaluation by link prediction does not rank the nonsensical triple (A Room With A View, director/film, James Ivory). Wang et al. [40] showed that models often assign high scores to such invalid triples. To assess how models rank all possible triples, even the nonsensical ones, for a given relation, we performed the entity-pair ranking evaluation protocol on the two datasets mentioned above (Section 4.2). Another limitation of link prediction is its assumption that an entity’s properties are already known, focusing only on retrieving the correct object or subject entities. In practice, determining whether a property applies to an entity is itself a challenge. To address this gap, we introduce a new evaluation protocol, property prediction (detailed in Section 4.3). A third alternative evaluation protocol to link prediction, triple classification [41, 35], assesses whether a given triple is a true or false fact. This requires negative triples in the test and validation sets. Traditionally, these negative triples are generated by randomly replacing a head or tail entity, an approach that does not sufficiently challenge the models [35, 41]. Inspired by hard negatives suggested by Pezeshkpour et al. [26] and Safavi et al. [31], we propose focusing on generating type-consistent negatives using the Freebase type system as models generally excel at identifying entity types, making type-inconsistent negatives easier to classify and providing less meaningful evaluation challenges (detailed in Section 4.4).
Prior research has investigated various aspects of KGE evaluation. Several studies have explored the impact of training strategies on KGE models’ performance [3]. Others examined biases in KGE benchmarks [32]. Additionally, some works highlighted the limitations and challenges associated with existing evaluation datasets [2, 34]. Evaluation metrics for KGE models have been extensively studied. For instance, Hoyt et al. [14] proposed desiderata for improved metrics to enhance interpretability and comparability of existing metrics across datasets of different sizes and properties. Similarly, Hubert et al. [17] proposed $Sem@K$, a semantic oriented metric, to evaluate the semantic-awareness of KGE models. Other studies explored different evaluation protocols [40, 30] and assessed KGE models from an information retrieval perspective [48]. Additionally, KGE models have been evaluated via real-world applications, such as recommendation systems [13]. ReliK [10] introduced a framework to quantify a KGE model’s reliability for specific downstream tasks without requiring their direct execution. Furthermore, some studies focused on what the embeddings learn and whether similar entities have similar embeddings [18].
@@
  • We highlight the limitations of link prediction as an evaluation protocol for KGE models, providing quantitative results underscoring the need for diverse evaluation protocols, and demonstrate how dataset size affects model performance and relative rankings among different KGE approaches.
@@
To perform the experiments in this paper, we utilized the provided training, validation, and test sets of the FB-CVT-REV and FB+CVT-REV datasets, with the split ratio of 90/5/5. Our experiments were conducted on an NVIDIA H100 80GB GPU. All the scripts and datasets used in our experiments, and additional details on the experimental setup are publicly available at https://github.com/idirlab/largeKGEeval.
@@ 2 Background
2.1 Most Common Evaluation Protocols and Evaluation Metrics
@@ 2.2 Evaluation Datasets and Frameworks
3 Inadequacy of Evaluation Metrics
@@
3.1 Closed-World Assumption
The aforementioned performance metrics are designed to evaluate a model’s performance based on the closed-world assumption. The concepts of closed-world and open-world assumptions were first introduced by Reiter [28, 27]. Suppose that the total knowledge of a domain is available. Then it is not necessary to explicitly represent negative facts; they are inferred from the absence of their positive counterparts. Reiter [28] called this the closed-world assumption. It can be stated as “if $P$ is not provable from the knowledge base, assume $\neg P$,” where P is a positive fact. This assumption is far from realistic due to the incompleteness of general real-world knowledge graphs and thus has flaws when a model correctly predicts a triple that does not exist in a dataset. More specifically, if a corrupted triple of a given test triple does not exist in the current dataset, it is considered false. However, the corrupted triple might actually be valid. If a model ranks the correct corrupted triple higher than the test triple itself, its accuracy measures will be penalized, which contradicts the primary goal of link prediction—finding correct triples that do not already exist in the dataset.
@@ 3.2 Global Metrics
Macro-averaging by relation.    Another limitation of the current metrics is that they aggregate the accuracy of all predicted triples into a single value—the micro-average accuracy—which masks the specific strengths and weaknesses of the models. The Freebase dataset exhibits a highly skewed distribution, with a few high-frequency relations and a large number of low-frequency ones [1]. The micro-average accuracy is predominantly influenced by the most frequent relations. To establish a metric that provides equal weight to all relations in the evaluation, we also propose incorporating the macro-average of per-relation accuracy. The macro-averaged MRRis defined as $\frac{1}{|R|}\sum_{r\in R}(\frac{1}{2\ |T_{r}|}\sum_{(\textsf{\small h},\textsl{\small r},\textsf{\small t})\in T_{r}}(\frac{1}{\emph{rank}_{\textsf{\small h}}}+\frac{1}{\emph{rank}_{\textsf{\small t}}}))$, where $|R|$ is the total number of distinct relations and $|T_{r}|$ is the size of the test triples in relation $r$. Table 2 compares micro- and macro-averaged MRR. As observed, MRR improves across all models with macro-averaging, indicating that treating all relations equally results in higher overall performance. This suggests that models perform relatively well even on low-frequency relations.
@@
Left vs. right entity link prediction.    Moreover, we propose evaluating left and right entity link prediction—(?, r, t) and (h, r, ?), respectively—separately for each relation, instead of relying solely on the overall average accuracy of test triples. Table 2 illustrates that models have considerably higher accuracy on right entity link predictions than left entity link predictions. This distinction reveals important asymmetries in model performance; thus, examining predictions in both directions separately provides a more nuanced understanding of model capabilities.
@@
N-ary vs. binary relations. In the presence of n-ary relationships (i.e., CVT nodes), link prediction becomes more challenging since these nodes are often long-tail entities with limited connectivity [34]. This limited connectivity stems from the fact that each CVT node typically represents the specific attributes or relationships of just a single main entity, inherently restricting its connections within the graph. However, their impact on the effectiveness of current link prediction models remains underexplored. To investigate this, we categorized relations into two types—binary relations, which connect two regular entities, and n-ary (or concatenated) relations. In FB+CVT-REV, n-ary relations link regular entities to CVT nodes, while in FB-CVT-REV concatenated relations represent binary conversions of n-ary relations. Table 3 presents the results for both relation types, with the “all” columns aggregating them. The results indicate that evaluation on concatenated relations yields overestimated accuracy metrics for most models compared to their performance on n-ary ones. This may result from decomposing n-ary relations into several binary relations through Cartesian product-like conversions. The decomposition allows models to more easily learn these simplified patterns when evaluated separately, creating an illusion of higher capability of the models.color=yellowcolor=yellowtodo: color=yellowTransE is different. are we able to explain it? Nasim: I am not sure how to explain this Moreover, across all models and datasets, the accuracy on concatenated/n-ary relations is significantly higher than binary relations because they exhibit more consistent structural patterns with lower variability, making them easier for models to learn despite their origin in complex relationships.
@@
Macro-averaging by domain.   Freebase is a cross-domain common fact KG that comprises factual information in a broad range of domains [34], such as music, sports, biology, etc. We analyzed link prediction performance on the 12 most frequent domains in FB-CVT-REV, which together account for nearly 95% of all subject-matter triples—those representing real-world facts, as opposed to metadata that describes the data itself [34]. Table 4 reports the number of distinct relations, triples, and MRR for each domain. The results show significant variance in models’ accuracy across domains, with overall micro-averaged accuracy heavily influenced by dominant domains, particularly music, which accounts for over 60% of the triples. This highlights the importance of incorporating macro-average evaluation alongside micro-average evaluation to give equal weight to each domain, a factor largely overlooked in the existing literature. Here, the macro-average is defined at the domain level rather than the relation level. The domain-level macro-averaged MRRis $\frac{1}{|D|}\sum_{d\in D}(\frac{1}{2\ |T_{d}|}\sum_{(\textsf{\small h},\textsl{\small r},\textsf{\small t})\in T_{d}}(\frac{1}{\emph{rank}_{\textsf{\small h}}}+\frac{1}{\emph{rank}_{\textsf{\small t}}}))$, where $|D|$ is the total number of distinct domains and $|T_{d}|$ is the size of the test triples in domain $d$. Given the large variance of performance across different domains, it is also important to understand the models’ performance on a domain is influenced by each domain’s characteristics. For instance, the models had almost perfect accuracy on domain measurement_unit. Most of the frequent relations in this domain are 1-to-n relations—relations in which a single head entity can be associated with multiple tail entities. One example is /measurement_unit/dated_integer/source-/location/hud_foreclosure_area/estimated_number_foreclosures, which is a concatenated property. For instance, when the subject is United States Department of Housing and Urban Development, the objects are Lake Barcroft, Van Vleck, Fort Ahby, and so on. Given the similarity between the objects, it could be easier to predict the subject. On the other hand, the models had particularly low accuracy on domains such as book. One of the most prevalent relations in this domain is /book/book_edition/isbn. An ISBN value mostly appears in one triple only, with highly rare exceptions, and that triple would be an instance of /book/book_edition/isbn. To predict either the book or the ISBN value, without the ISBN value being connected to anything else, the model would be helpless.

### v2

@@ Abstract
Knowledge graph embedding (KGE) models are extensively studied for knowledge graph completion, yet their evaluation remains constrained by unrealistic benchmarks. Standard evaluation metrics rely on the closed-world assumption, which penalizes models for correctly predicting missing triples, contradicting the fundamental goals of link prediction. These metrics often compress accuracy assessment into a single value, obscuring models’ specific strengths and weaknesses. The prevailing evaluation protocol, link prediction, operates under the unrealistic assumption that an entity’s properties, for which values are to be predicted, are known in advance. While alternative protocols such as property prediction, entity-pair ranking, and triple classification address some of these limitations, they remain underutilized. Moreover, commonly used datasets are either faulty or too small to reflect real-world data. Few studies examine the role of mediator nodes, which are essential for modeling n-ary relationships, or investigate model performance variation across domains. This paper conducts a comprehensive evaluation of four representative KGE models on large-scale datasets FB-CVT-REV and FB+CVT-REV. Our analysis reveals critical insights, including substantial performance variations between small and large datasets, both in relative rankings and absolute metrics, systematic overestimation of model capabilities when n-ary relations are binarized, and fundamental limitations in current evaluation protocols and metrics.
@@ 1 Introduction
This paper presents a comprehensive large-scale evaluation of KGE models for KG completion. Although these models have been extensively evaluated to assess their efficacy and applicability in real-world applications, we demonstrate fundamental limitations in current evaluation methodologies that result in inaccurate assessments of model performance. Our investigation focuses particularly on large-scale analysis, examining two main interrelated issues: 1) shortcomings of standard rank-based evaluation metrics and 2) inherent weaknesses in evaluation protocols. Our findings reveal that these methodological limitations substantially affect how KGE models are understood and compared. We also examine how data representation choices affect performance outcomes by using two datasets that employ different data modeling approaches in all our experiments.
@@
Many existing evaluations of KGE models rely on datasets, such as FB15k-237 [6] and WN18RR [9], both of which exhibit several limitations. First, they fail to capture n-ary relationships among entities, which are modeled using mediator nodes (also called CVT nodes) in the Freebase data dump [25] and statement nodes in Wikidata. Our earlier findings [34] indicate that the conversion of n-ary relationships to binary ones, as implemented in FB15k-237, leads to substantial overestimation of KGE models’ performance. Second, these datasets are considerably smaller than real-world datasets. Experiment results from previous studies [34, 47] reveal significant differences in the performance of models between small-scale and large-scale datasets, with relative performance rankings among models also varying based on dataset size. Our experiments are conducted on two large-scale datasets, FB-CVT-REV and FB+CVT-REV [34]. While FB+CVT-REV represents n-ary relationships using CVT nodes, FB-CVT-REV converts n-ary relationships into binary ones, similar to FB15k-237. This distinction allows us to assess how KGE model performance varies when data is modeled differently. These datasets, derived from Freebase, are large-scale, represent real-world scenarios, and do not include the data redundancy that exists in FB15k. A discussion of these two datasets and other benchmarks is provided in Section 3.
@@
(1) Shortcomings of standard rank-based evaluation metrics: KG completion methods aim to identify missing but correct triples in a KG. However, current evaluation metrics paradoxically penalize models for correctly predicting triples that are missing from the evaluation dataset. Standard ranking-based metrics such as Mean Rank (MR) and Mean Reciprocal Rank (MRR) [45] operate under the closed-world assumption, treating all predictions outside the KG as incorrect. This contradicts the open-world assumption, which acknowledges that absent facts are not necessarily false [28]. Furthermore, each of these metrics aggregates performance across all relations and triples into a single value, obscuring models’ strengths and weaknesses on different relations. In Section 4, we explore the shortcomings of existing evaluation metrics with supporting experimental evidence discussed in Section 5. We illustrate how these shortcomings impact models’ performance and propose refinements to improve evaluation reliability.
@@
(2) Inherent weaknesses in evaluation protocols: The standard evaluation protocol for KGE models, link prediction [6], evaluates models by predicting the missing h in triple (?, r, t) or missing t in (h, r, ?). It assesses whether a model prioritizes correct answers over incorrect ones. However, as Wang et al. [40] pointed out, this approach can be misleading because it does not verify whether correct triples are ranked better than all type of false and nonsensical ones. For example, for a test triple (James Ivory, director/film, A Room With A View), evaluation by link prediction does not rank the nonsensical triple (A Room With A View, director/film, James Ivory). Wang et al. [40] showed that models often assign high scores to such invalid triples. To assess how models rank all possible triples, even the nonsensical ones, for a given relation, we performed the entity-pair ranking evaluation protocol on the two datasets mentioned above. Another limitation of link prediction is its assumption that an entity’s properties are already known, focusing only on retrieving the correct object or subject entities. In practice, determining whether a property applies to an entity is itself a challenge. To address this gap, we introduce a new evaluation protocol, property prediction. A third alternative evaluation protocol to link prediction, triple classification [41, 35], assesses whether a given triple is a true or false fact. This requires negative triples in the test and validation sets. Traditionally, these negative triples are generated by randomly replacing a head or tail entity, an approach that does not sufficiently challenge the models [35, 41]. Inspired by hard negatives suggested by Pezeshkpour et al. [26] and Safavi et al. [31], we propose focusing on generating type-consistent negatives using the Freebase type system as models generally excel at identifying entity types, making type-inconsistent negatives easier to classify and providing less meaningful evaluation challenges. Further details on the limitations of link prediction can be found in Section 4, while experiments involving alternative evaluation protocols are presented in Section 5.
@@
  • We highlight the limitations of link prediction as an evaluation protocol for KGE models, and demonstrate how dataset size affects model performance and relative rankings among different KGE approaches.
@@
All the scripts and datasets used in our experiments, and additional details on the experimental setup, are publicly available at https://github.com/idirlab/largeKGEeval.
2 Related work
Prior research has investigated various aspects of KGE evaluation. Several studies have explored the impact of training strategies on KGE models’ performance [3]. Evaluation metrics for KGE models have been extensively studied. For instance, Hoyt et al. [14] proposed desiderata for improved metrics to enhance interpretability and comparability of existing metrics across datasets of different sizes and properties. Similarly, Hubert et al. [17] proposed $Sem@K$, a semantic-oriented metric, to evaluate the semantic-awareness of KGE models. Other studies explored different evaluation protocols [40, 30] and assessed KGE models from an information retrieval perspective [48]. Additionally, KGE models have been evaluated via real-world applications, such as recommendation systems [13]. ReliK [10] introduced a framework to quantify a KGE model’s reliability for specific downstream tasks without requiring their direct execution. Furthermore, some studies focused on what the embeddings learn and whether similar entities have similar embeddings [18]. Others examined biases in KGE benchmarks [32, 2]. The most closely related work to ours is [34], which investigates the impact of various data modeling idiosyncrasies, such as representing n-ary relations via mediator nodes, on the performance of KGE models at scale. Additionally, they provide several well-processed datasets that facilitate more realistic assessments of KGE models, two of which are used in our experiments. In this paper, we primarily focus on the limitations of evaluation metrics and protocols, rather than solely on data modeling. Nevertheless, all our experiments place emphasis on large-scale evaluation and systematically examine the effect of different data modeling choices on model performance as well.
@@ 3 Background
3.1 Common Evaluation Protocols and Evaluation Metrics
@@ 3.2 Evaluation Datasets and Frameworks
4 Limitations of Current Evaluation Methodologies
4.1 Shortcomings of Standard Evaluation Metrics
4.1.1 Closed-World Assumption
The aforementioned performance metrics are designed to evaluate a model’s performance based on the closed-world assumption. Suppose that the total knowledge of a domain is available. Then it is not necessary to explicitly represent negative facts; they are inferred from the absence of their positive counterparts. This concept of closed-world assumption, first introduced by Reiter [28, 27], can be stated as “if a positive fact $P$ is not provable from the knowledge base, assume $\neg P$.” This assumption is far from realistic due to the incompleteness of real-world knowledge graphs and thus has flaws when a model correctly predicts a triple that does not exist in a dataset. More specifically, if a corrupted triple of a given test triple does not exist in the current dataset, it is considered false. However, the corrupted triple might actually be valid. If a model ranks the correct corrupted triple higher than the test triple itself, its accuracy measures will be penalized, which contradicts the primary goal of link prediction—finding correct triples that do not already exist in the dataset.
4.1.2 Global Metrics
Another limitation of the current metrics is that they aggregate the accuracy of all predicted triples into a single value—the micro-average accuracy—which masks the specific strengths and weaknesses of the models. The Freebase dataset exhibits a highly skewed distribution, with a few high-frequency relations and a large number of low-frequency ones [1]. Thus, the micro-average accuracy is predominantly influenced by the most frequent relations. Moreover, the practice of reporting a single, overall accuracy neglects important sources of variability, such as asymmetries between predicting head versus tail entities and performance variation across semantic domains, which can all affect a model’s real-world utility. Furthermore, in the presence of n-ary relationships (i.e., CVT nodes), link prediction becomes more challenging since these nodes are often long-tail entities with limited connectivity [34]. This limited connectivity stems from the fact that each CVT node typically represents the specific attributes or relationships of just a single main entity, inherently restricting its connections within the graph. However, their impact on the effectiveness of current link prediction models remains underexplored.
4.2 Inherent Weaknesses in Evaluation Protocols
Link prediction is a widely used evaluation protocol for assessing KGE models. However, it has certain limitations. First, its evaluation setup can be misleading as we cannot verify if a model ranks all types of false or nonsensical triples lower than correct triples. For example, consider a test triple (James Ivory, director/film, A Room With A View). Evaluation by link prediction only seeks responses to two sensible questions: “who directed the film A Room With A View?”, i.e., (?, director/film, A Room With A View), and “which film(s) are directed by James Ivory?”, i.e., (James Ivory, director/film, ?). It does not verify if a model would rank false or nonsensical triples, such as (A Room With A View, director/film, James Ivory) lower than the correct one. In this triple, the subject and object entities are swapped, resulting in an incorrect direction for the relation. Wang et al. [40] discussed this problem and proposed to use an alternative protocol called entity-pair ranking. Second, link prediction is based on the assumption that the presence of a particular property on an entity is already known. The evaluation focuses on whether a model can derive the correct property values. In reality, though, it remains a challenge to determine whether a property is valid for a given entity in the first place. Therefore, we propose the property prediction task as another evaluation protocol. Both entity-pair ranking and property prediction, along with a third alternative protocol, triple classification, are discussed in Section 5.2.
5 Experiments
In conducting our experiments, we utilized the provided training, validation, and test sets of the FB-CVT-REV and FB+CVT-REV datasets, with the split ratio of 90/5/5. Our experiments were performed on an NVIDIA H100 80GB GPU.
5.1 Empirical Analysis and Improvements of Evaluation Metrics
@@
5.1.1 Closed-World Assumption
@@ 5.1.2 Macro-Averaging by Relation
To establish a metric that provides equal weight to all relations in the evaluation, we propose incorporating the macro-average of per-relation accuracy. The macro-averaged MRRis defined as $\frac{1}{|R|}\sum_{r\in R}(\frac{1}{2\ |T_{r}|}\sum_{(\textsf{\small h},\textsl{\small r},\textsf{\small t})\in T_{r}}(\frac{1}{\emph{rank}_{\textsf{\small h}}}+\frac{1}{\emph{rank}_{\textsf{\small t}}}))$, where $|R|$ is the total number of distinct relations and $|T_{r}|$ is the size of the test triples in relation $r$. Table 2 compares micro- and macro-averaged MRR. As observed, MRR improves across all models with macro-averaging, indicating that treating all relations equally results in higher overall performance. This suggests that models perform relatively well even on low-frequency relations.
5.1.3 Left vs. Right Entity Link Prediction
@@
We propose evaluating left and right entity link prediction—(?, r, t) and (h, r, ?), respectively—separately for each relation, instead of relying solely on the overall average accuracy of test triples. Table 2 shows that models have considerably higher accuracy on right entity link predictions than left entity link predictions. This distinction reveals important asymmetries in model performance; thus, examining predictions in both directions separately provides a more nuanced understanding of model capabilities.
5.1.4 N-ary vs. Binary Relations
@@
To investigate the impact of n-ary relationships on link prediction, we categorized relations into two types—binary relations, which connect two regular entities, and n-ary (or concatenated) relations. In FB+CVT-REV, n-ary relations link regular entities to CVT nodes, while in FB-CVT-REV concatenated relations represent binary conversions of n-ary relations. Table 3 presents the results for both relation types, with the “all” columns aggregating them. The results indicate that evaluation on concatenated relations yields overestimated accuracy metrics for most models compared to their performance on n-ary ones. This may result from decomposing n-ary relations into several binary relations through Cartesian product-like conversions. The decomposition allows models to more easily learn these simplified patterns when evaluated separately, creating an illusion of higher capability of the models. Moreover, across all models and datasets, the accuracy on concatenated/n-ary relations is significantly higher than binary relations because they exhibit more consistent structural patterns with lower variability, making them easier for models to learn despite their origin in complex relationships.
@@
5.1.5 Macro-Averaging by Domain
Freebase is a cross-domain common fact KG that comprises factual information in a broad range of domains [34], such as music, sports, biology, etc. We analyzed link prediction performance on the 12 most frequent domains in FB-CVT-REV, which together account for nearly 95% of all subject-matter triples—those representing real-world facts, as opposed to metadata that describes the data itself [34]. Table 4 reports the number of distinct relations, triples, and MRR for each domain. The results show significant variance in models’ accuracy across domains, with overall micro-averaged accuracy heavily influenced by dominant domains, particularly music, which accounts for over 60% of the triples. This highlights the importance of incorporating macro-average evaluation alongside micro-average evaluation to give equal weight to each domain, a factor largely overlooked in the existing literature. Here, the macro-average is defined at the domain level rather than the relation level. The domain-level macro-averaged MRRis $\frac{1}{|D|}\sum_{d\in D}(\frac{1}{2\ |T_{d}|}\sum_{(\textsf{\small h},\textsl{\small r},\textsf{\small t})\in T_{d}}(\frac{1}{\emph{rank}_{\textsf{\small h}}}+\frac{1}{\emph{rank}_{\textsf{\small t}}}))$, where $|D|$ is the total number of distinct domains and $|T_{d}|$ is the size of the test triples in domain $d$. Given the large variance of performance across different domains, it is also important to understand the models’ performance on a domain is influenced by each domain’s characteristics. For instance, the models had almost perfect accuracy on domain measurement_unit. Most of the frequent relations in this domain are 1-to-n relations—relations in which a single head entity can be associated with multiple tail entities. One example is /measurement_unit/dated_integer/source-/location/hud_foreclosure_area/estimated_number_foreclosures, which is a concatenated property. For instance, when the subject is United States Department of Housing and Urban Development, the objects are Lake Barcroft, Van Vleck, Fort Ahby, and so on. Given the similarity between the objects, it could be easier to predict the subject. On the other hand, the models had particularly low accuracy on domains such as book. One of the most prevalent relations in this domain is /book/book_edition/isbn. An ISBN value mostly appears in one triple only, with highly rare exceptions, and that triple would be an instance of /book/book_edition/isbn. To predict either the book or the ISBN value, without the ISBN value being connected to anything else, the model would be helpless.